Circular No. 7515

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 U5 (LINEAR) An object with unusual motion that was reported as asteroidal by

the LINEAR survey (discovery observation below) and posted on the NEO Confirmation Page has been found to be cometary by other observers. The object seemed diffuse with a 14′′ coma and 18′′ tail in p.a. 170^{o} on CCD images obtained by J. Tichá and M. Tichý (Kleť) on Oct. 30.0 UT. Images taken on Oct. 30.2 by D. A. Klinglesmith III (Socorro, NM) show slight diffuseness and a tail ∼ 20′′ long in p.a. 220^{o}; images by Y. Ikari (Moriyama, Japan) also show a tail in p.a. 220^{o} on Oct. 30.6. J. Biggs (Perth Observatory) notes that images of C/2000 U5 were larger than nearby stars and elongated toward the south-southwest on Oct. 31.6. D. T. Durig (Sewanee, TN) found a tail ∼ 25′′ long in p.a. 170^{o} on Nov. 1.4 images.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Oct. 29.38178  5^{h}06^{m}33.^{s}49  +2^{o}04′56′′.6  17.1

MPEC 2000-V02 contains the available astrometry and the parabolic orbital elements that produced the following ephemeris:

2000TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 23 5^{h}13.^{m}36 + 1^{o}26.^{′}7 2.699 3.413 128.^{o}9 13.^{o}1 14.5
Nov. 2 5 02.36 + 2 28.2 2.679 3.501 140.4 10.4 14.6
12 4 49.81 + 3 38.0 2.687 3.590 151.8  7.5 14.7
22 4 36.45 + 4 55.0 2.729 3.678 161.2  5.0 14.8
Dec. 2 4 23.16 + 6 17.3 2.809 3.767 164.1  4.1 15.0

COMET P/2000 U6 (TICHÝ) An object found by Milos Tichý on images taken at Kleť with J. Tichá

and M. Kočer on Oct. 23 (discovery observation below), originally reported as asteroidal, was subsequently noted to be diffuse on Kleť images taken during Oct. 28.9–29.2 UT; Tichý also noted a 10′′ coma on Oct. 29.8 images. Images obtained on Nov. 1 by S. Sanchez and M. Blasco at Mallorca and by D. T. Durig at Sewanee also showed diffuseness. MPEC 2000-V03 contains the available astrometry and orbital elements (T = 2000 Oct. 4.6 TT, i = 19^{o}.3, q = 2.150 AU, P = 7.3 yr).

2000 UT  α_{2000}  δ_{2000}  m_{2}
Oct. 23.07698  3^{h}06^{m}57.^{s}01  +28^{o}51′38′′.9  17.8
2000 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green